OpenStudy (anonymous):

The price of a large pizza increased from $10 to $12. What was the percent of increase in the cost of the pizza? Round to the nearest percent, if necessary. A. 17% B. 20% C. 2% D. 83%

OpenStudy (anonymous):

2 is what percent of 10?

OpenStudy (anonymous):

5%

OpenStudy (anonymous):

@zimmah

OpenStudy (anonymous):

\[\large \frac{ 2 }{ 10 }=\frac{ x }{ 100 }\]

OpenStudy (anonymous):

what is x?

OpenStudy (anonymous):

50..?

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

how did you arrive at 50?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

what is 2 divided by 10?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

actually, that's 10 divided by 2

OpenStudy (anonymous):

whats the difference??

OpenStudy (anonymous):

well, the same difference as for example 5-3=2 and 3-5=-2. The order of division is important

OpenStudy (anonymous):

If you divide a number by 10, you simply move the decimal one place to the left. So 2 becomes 0.2 and 20 becomes 2.0, and so on.

OpenStudy (anonymous):

2 divided by 10 is 0.2 and 10 divided by 2 is 5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it 2%

OpenStudy (anonymous):

close, but not quite, you'll have to multiple 0.2 by 100 to get the percentage

OpenStudy (anonymous):

20%

OpenStudy (anonymous):

a percentage is a 1/100th part of something.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Clare has 130 songs in her music collection. Fifty-nine of the songs are rock and roll. About what percent of the songs are rock and roll? A. 41% B. 45% C. 59% D. 71%

OpenStudy (anonymous):

because if the pizza was $100 to begin with, and the pizza would increase with the same percentage, the pizza would be $120 an increase of $20 per $100, or 20%

OpenStudy (anonymous):

oops my bad

OpenStudy (anonymous):

\[\large \frac{ 59 }{ 130 }*100\]

OpenStudy (anonymous):

i got 45% is that right..?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what is the solution to the proportion 9/X= 4/12

OpenStudy (anonymous):

start by multiplying both sides with 12

OpenStudy (anonymous):

is it 17..?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh wait multiplying by 12 may not be the best thing to do

OpenStudy (anonymous):

okay i got 9 times 12=108 and 4 times 12= 48

OpenStudy (anonymous):

\[\large \frac{ 9 }{ X }=\frac{ 4 }{ 12 }\] flip both sides\[\large \frac{ X }{ 9 }=\frac{ 12 }{ 4 }\]

OpenStudy (anonymous):

yes i figured that would only make it more confusing, sorry for that.

OpenStudy (anonymous):

27..?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

A bag contains 8 oatmeal cookies and 10 molasses cookies. What is the ratio, written in lowest terms, of oatmeal cookies to the total number of cookies in the bag? A. 5 to 4 B. 4 to 5 C. 5 to 9 D. 4 to 9

OpenStudy (anonymous):

there's 18 cookies in the bag, and 8 of them are oatmeal

OpenStudy (anonymous):

so \[\frac{ 8 }{ 18 }\]

OpenStudy (anonymous):

which can be simplified to ....

OpenStudy (anonymous):

4/9..?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Three tickets to a concert cost $54. How much do 7 tickets cost? A. $58 B. $72 C. $108 D. $126

OpenStudy (anonymous):

i get 135 every time

OpenStudy (anonymous):

close, but it's 126. $54*7/3